# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

### v2

@@ 4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
Participants imagined that in addition to private homes and personal spaces, SMH neighborhoods would include shared third places where people could casually gather, lowering social barriers and creating opportunities for connection. P05 described these areas as “neutral spaces”, not tied to any specific activity or relationship status, where groups could gather for a range of purposes. They imagined “some places to sit, and some kind of flat table-like surface” that could support casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces, they explained, would function as “a third, neutral territory” where people could interact more organically, without the friction of managing invitations or switching platforms.
Participants described shared social spaces within neighborhoods, such as common rooms, courtyards, or clubhouses, where residents could casually interact outside their personal spaces. P01 described leaving one’s home to gather in a “courtyard or game room” at the center of the neighborhood. Drawing on examples from Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms that cater to different interests, such as a Quidditch field or library, creating opportunities for people to meet others through shared activities. P09 similarly imagined a “virtual dining room” that would allow people to casually talk while doing something familiar, like eating together. Unlike the more segmented house-based accommodation at Hogwarts, this dining space was described as open and inclusive, allowing people to connect across social groups.
Others described third places located in public areas that supported casual, serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” Similarly, P08 suggested a marketplace space where selling items could spark conversations: “‘Hey, look! This is for sale’… have a talk about it… and meet new people.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you…and meet new people.” Other suggestions included “public study hall[s]” (P17) where participants could “interact with everybody,” (P17) leading them to experience that, “all of a sudden, you know this person that plays the piano.” (P17)
@@ 4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized the importance of filtering out irrelevant content, such as celebrity drama or political commentary, and prioritizing updates from close friends and trusted communities. P19 criticized how platforms like Instagram have shifted away from their initial focus:
@@
Several participants advocated minimizing celebrity influence. P02 stated, “I don’t want celebrity stuff… I don’t want any celebrity drama,” while P01 suggested removing celebrities altogether if the platform prioritized personal enjoyment. If included, they proposed isolating them to a “celebrity zone.” Prioritizing posts from close friends was a key preference. P07 explained, “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH offering greater control over interactions and content, noting it would be “more designed for people you actually know… like your friends, classmates, or Quidditch team,” rather than celebrities or acquaintances. P19, similarly, envisioned that on SMH, they could “filter out for friends and family, and even then condense your circle a little bit more.”
@@ 4.3.4. Developing Socially Grounded and Themed Conversational Spaces
Participants described how themed, socially grounded conversational spaces designed for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
@@
Participants described how the design and feel of conversational spaces could make social interactions on SMH feel safer, more personal, and more inviting than on existing platforms. Many valued the idea of walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to a broad, undefined audience. P22, for example, imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences to help spark meaningful conversations. They described these as “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting them with the more exposed, contextless feel of mainstream social media where “you don’t know everything around it.”
@@ 4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
To ensure SMH feels like a safe, personal environment, participants stressed the need for more comprehensive boundary controls beyond privacy tools found on existing platforms that they perceived as limited and inconsistent. P22, for example, explained that even with contact syncing turned off on Instagram, the platform still suggested their Finsta (Fake + Instagram; secondary account (Huang and Vitak, 2022)) account to people they were trying to avoid, based on mutual follow relationships. They stressed the need for stronger, more thorough blocking controls, including the ability to block “alternate accounts” and fully remove any trace of a blocked person’s presence, explaining, “I don’t wanna be able to tell that the person I blocked is even around me.” P14 similarly pointed out that even private accounts on Instagram still expose personal content, like Story Highlights, to “a random person”, revealing perceived gaps in platform-level privacy protections (Bojkov, 2019).
Participants expressed a desire for stronger control over their own visibility, imagining features that would let them opt out of being seen or discovered entirely. P03 suggested a feature like an “invisibility cloak” for moments when users “don’t want anyone to see you.” P07 proposed a “super private” profile mode where users “can’t even be searched… like you’re not even on the app” unless they choose to make themselves visible.
Preventing unsolicited contact from strangers was another recurring priority. P22 proposed a model inspired by Discord, where only people you’ve already interacted with can contact you, and others “can’t even be discovered unless you have something in common.” This idea resonated with P13, who argued that new connections should require mutual consent, suggesting that people “should only be able to add you if you meet in person and exchange a QR code.” They shared a personal experience of being able to find a stranger online based on just their first name, reflecting that “it’s too easy to find people on social media.”
@@ 4.4. Nuanced Self-Presentation
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This carefully curated image culture creates “a lot of seriousness” (P23), which is reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared an anecdote about helping a friend prepare an Instagram post, including selecting the “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting the full range of who they are, without the pressures of curation or performance.
@@ 4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
Participants consistently described customizable avatars and personal spaces as meaningful tools for self-expression, offering more creativity, personality, and authenticity than the static profiles or limited settings found on MSM. Unlike MSM, which typically offers only “two modes” (P10), i.e., “light and dark” (P10), participants imagined a much richer and more immersive experience. For example, P21 envisioned a fully customizable world where users could build their own “mini world” within a larger virtual space, complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school.” These spaces could be shared with friends as a way of “showing a little part of yourself.” Similarly, P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture,” to reflect their personal interests.
@@
Customizable “3D avatar[s]” (P11) and profiles were also seen as key ways of showing individuality. P22 reflected on the early days of Myspace (mys, [n. d.]), describing how they enjoyed coding their profile layouts and changing “the font and really make it reflect me.” Even small details, like choosing a username, were described as ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted that such forms of expression are often “mainly for you,” not just for public display. P17 added that even if users don’t consciously intend to reveal something personal, their choices still do: “your brain conjured that up, right? So I think that’s part of you.”
@@
Beyond visual customization, participants imagined SMH supporting more ambient, indirect forms of expression, offering subtle cues about what someone is doing, feeling, or thinking without needing to make a formal post. For instance, P01 described wanting their avatar to reflect real-time states, like the outfit they are wearing or the song they are listening to, so that friends could intuitively notice things like, “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ Similarly, P10 imagined that friends could pick up on changes in their interests simply by observing how they update their virtual space over time, providing “an insight into… what I’m thinking.”
@@
Several participants highlighted the contrast between the flat, compressed feel of 2D platforms and the immersive qualities of 3D environments. P10 noted that in 3D, others can see “as much stuff as you can see in 3D,” making it feel richer and more complete. P16 added that VR creates a sense of presence that feels more like meeting “real people… compared to the profile pictures,” because you can pick up on their “mannerisms” and notice details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
@@
Participants also described how these kinds of choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests. As P01 put it, such “creative expression” feels “a lot more personal” because it reflects “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without needing to start a direct conversation. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
@@ 4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides of themselves across various spaces and relationships, and reducing the pressure to maintain a single, polished identity. Rather than being limited to a single, polished identity, as is often the case on MSM, participants wanted ways to represent the full range of their interests, identities, and relationships. P01 described this as having a house with rooms “dedicated to each of your interests,” explaining, “I’m into multiple things. You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” This ability to represent different facets of oneself was seen as a way to reduce misjudgment and give others a more complete picture of who they are. P02 built on this idea, imagining that users with multiple “Muggle personalities” could organize them into separate spaces, saying, “Each personality has its own vault.”
@@
4.4.3. Sharing Everyday Activities and Lived Experiences to Deepend Understanding and Connection
@@
Participants valued the ability to share casual, everyday moments and rich, embodied memories, filling the gaps between the highlights usually shared on MSM into a more natural, meaningful extension of daily life and personal connection. P04 described wanting to share small, ordinary moments, like “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
@@
Participants also imagined SMH supporting the sharing of more personal, embodied memories as a way to deepen understanding and reduce misinterpretation. P04 described wanting to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pen, [n. d.]) as a way to create stronger emotional connections. Unlike MSM, where communication often feels like “just words,” (P03) participants saw memory-sharing as a way to communicate context, emotion, and intention more fully. P02 explained that this kind of sharing could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
@@
Some participants imagined that these shared memories could include sensory or emotional details, not just visuals. P01, for example, expressed interest in being able to “transmit smells through social media,” adding another layer to digital communication. Others compared it to live streaming personal experiences. P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness in capturing fleeting moments, like crossing a state line, without the need to pause and record. P03 imagined being able to instantly share moments with loved ones “in the moment,” rather than having to explain them after the fact, when the emotion or relevance might already have passed.
@@ 4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
Participants emphasized that flexible, contextual privacy is essential for supporting authentic self-presentation, allowing people to manage who sees what, when, and how, without feeling overexposed or misrepresented. They described current platforms as offering rigid and limited controls that fail to account for the complexity of real social life. P16 explained that on Instagram, for example, privacy options are reduced to features like “close friends,” adding, “it’s kind of a pain to organize.” P11 similarly noted that as people share “more like their genuine real life and their experiences,” privacy needs to become “a little bit more complicated,” requiring finer-grained control over visibility.
@@ 4.5. Intentional Engagement
Participants described their growing frustration with how current social media platforms prioritize passive consumption over meaningful connection. In response, they envisioned SMH as a space that promotes intentional, socially rewarding interactions. This section explores four key themes that reflect their desire to refocus social media on authentic connection, purposeful engagement, and meaningful use of time.
@@ 4.5.1. Encouraging Deliberate Navigation and Content Discovery
Participants imagined SMH as a space that encourages deliberate, effortful discovery, replacing the passive, algorithm-driven feeds of MSM with experiences that require users to actively choose what and who they engage with. P22 described their frustration with Instagram, explaining that “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage more intentional, purposeful engagement, both when browsing content and when navigating the virtual world.
@@
Participants also imagined features that would prompt users to actively choose what they wanted to engage with before entering a space. P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces that would surface content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
Physical movement through SMH’s world was also seen as key to this intentionality. Participants valued having to walk, explore, and navigate to find people, spaces, or activities, turning even small actions into social moments of presence and care. P03 explained that this kind of embodied movement makes SMH feel “more connected than regular social media,” (P03) because it requires active participation rather than passive scrolling.
This desire for intentional interaction extended to how users move through the platform itself. Participants imagined physical navigation, such as walking or traveling through virtual space, as a way to make engagement feel more meaningful. P02 explained that they would prefer slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.” While features like “spells” (P06) or the “Marauder’s Map” (P21) were imagined as convenient options, participants valued the ability to walk and explore, turning simple actions like visiting a friend or browsing a shop into shared social moments. P03 explained that this kind of embodied interaction feels “more connected than regular social media” (P03) because it requires active presence rather than passive scrolling.
@@
Participants believed that this kind of effortful navigation would naturally discourage casual lurking or influencer-driven engagement, shifting the platform’s focus toward genuine connection. As P05 put it, “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”

### v3

@@ 4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
Participants imagined SMH neighborhoods including shared third places where people could casually gather, lowering social barriers and creating connection opportunities. P05 described these as “neutral spaces” not tied to specific activities or relationships, where groups could gather for various purposes. They imagined “some places to sit, and some kind of flat table-like surface” supporting casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces would function as “a third, neutral territory” for organic interaction without managing invitations or switching platforms.
Participants described shared neighborhood spaces like common rooms, courtyards, or clubhouses where residents could casually interact outside personal spaces. P01 described gathering in a “courtyard or game room” at the neighborhood center. Drawing on Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms catering to different interests, like a Quidditch field or library, creating opportunities to meet others through shared activities. P09 imagined a “virtual dining room” allowing casual conversation while eating together. Unlike Hogwarts’ segmented house-based accommodation, this dining space was open and inclusive, enabling cross-social group connections.
Others described public third places supporting casual, serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you.”
@@ 4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized filtering irrelevant content like celebrity drama or political commentary while prioritizing updates from close friends and trusted communities. P19 criticized how platforms like Instagram have shifted away from their initial focus:
@@
Several participants advocated minimizing celebrity influence, with P01 suggesting isolating them to a “celebrity zone.” P07 captured the preference for close connections: “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH being “more designed for people you actually know… like your friends, classmates, or Quidditch team.”
@@ 4.3.4. Developing Socially Grounded and Themed Conversational Spaces
Participants described how themed, socially grounded conversational spaces for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
@@
Participants described how conversational space design could make SMH interactions feel safer, more personal, and more inviting than existing platforms. Many valued walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to broad, undefined audiences. P22, for example, imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences to help spark meaningful conversations. They described these as “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting them with the more exposed, contextless feel of mainstream social media where “you don’t know everything around it.”
@@ 4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
To ensure SMH feels safe and personal, participants stressed needing comprehensive boundary controls beyond existing platforms’ limited and inconsistent privacy tools. P22 explained that even with contact syncing off, Instagram still suggested their Finsta (secondary account (Huang and Vitak, 2022)) to people they wanted to avoid, based on mutual follows. They stressed needing stronger, more thorough blocking controls, including blocking “alternate accounts” and fully removing blocked person traces, explaining “I don’t wanna be able to tell that the person I blocked is even around me.” P14 pointed out that even private Instagram accounts expose personal content like Story Highlights to “a random person”, revealing perceived privacy protection gaps (Bojkov, 2019).
Participants wanted stronger control over their visibility, imagining features letting them opt out of being seen or discovered entirely. P03 suggested an “invisibility cloak” for moments when users “don’t want anyone to see you.”
Preventing unsolicited stranger contact was another priority. P22 proposed a Discord-inspired model where only people you’ve interacted with can contact you, while P13 argued people “should only be able to add you if you meet in person and exchange a QR code,” reflecting that “it’s too easy to find people on social media.”
@@ 4.4. Nuanced Self-Presentation
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This curated image culture creates “a lot of seriousness” (P23), reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared helping a friend prepare an Instagram post, including selecting “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting their full range without curation or performance pressures.
@@ 4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
Participants consistently described customizable avatars and personal spaces as meaningful self-expression tools, offering far more creativity than MSM’s limited settings. P21 envisioned building their own “mini world” complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school,” sharing these spaces as “showing a little part of yourself.” P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture.”
@@
Customizable “3D avatar[s]” (P11) and profiles were key ways of showing individuality. P22 reflected on early Myspace (mys, [n. d.]), describing how they enjoyed coding profile layouts and changing “the font and really make it reflect me.” Even small details like choosing usernames were ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted such expression is often “mainly for you,” not just for public display. P17 added that even if users don’t consciously intend to reveal something personal, their choices still do: “your brain conjured that up, right? So I think that’s part of you.”
@@
Beyond visual customization, participants imagined SMH supporting ambient, indirect expression, offering subtle cues about what someone is doing, feeling, or thinking without formal posts. P01 wanted their avatar to reflect real-time states like current outfit or music, so friends could intuitively notice “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 imagined friends picking up on interest changes by observing virtual space updates over time, providing “an insight into… what I’m thinking.”
@@
Several participants highlighted the contrast between 2D platforms’ flat, compressed feel and 3D environments’ immersive qualities. P16 noted that VR creates presence feeling more like meeting “real people… compared to profile pictures,” because you can pick up on “mannerisms” and details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
@@
Participants described how these choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests. As P01 put it, such “creative expression” feels “a lot more personal” because it reflects “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without needing to start a direct conversation. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
@@ 4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides across various spaces and relationships while reducing pressure to maintain a single, polished identity. Rather than being limited to one polished identity, as on MSM, participants wanted ways to represent their full range of interests, identities, and relationships. P01 described this as having a house with rooms “dedicated to each of your interests,” explaining, “I’m into multiple things. You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” This ability to represent different facets of oneself was seen as a way to reduce misjudgment and give others a more complete picture of who they are. P02 built on this idea, imagining that users with multiple “Muggle personalities” could organize them into separate spaces, saying, “Each personality has its own vault.”
@@
4.4.3. Sharing Everyday Activities and Lived Experiences to Deepen Understanding and Connection
@@
Participants valued sharing casual, everyday moments and rich, embodied memories, filling gaps between MSM’s typical highlights to create a more natural, meaningful extension of daily life and personal connection. P04 described wanting to share small, ordinary moments, like “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
@@
Participants imagined SMH supporting personal, embodied memory-sharing to deepen understanding and reduce misinterpretation. P04 described wanting to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pen, [n. d.]) as a way to create stronger emotional connections. Unlike MSM, where communication often feels like “just words,” (P03) participants saw memory-sharing as a way to communicate context, emotion, and intention more fully. P02 explained that this kind of sharing could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
@@
Some participants imagined these shared memories including sensory or emotional details beyond visuals. P01, for example, expressed interest in being able to “transmit smells through social media,” adding another layer to digital communication. Others compared it to live streaming personal experiences. P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness in capturing fleeting moments, like crossing a state line, without the need to pause and record. P03 imagined being able to instantly share moments with loved ones “in the moment,” rather than having to explain them after the fact, when the emotion or relevance might already have passed.
@@ 4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
Participants emphasized that flexible, contextual privacy is essential for authentic self-presentation, allowing people to manage who sees what, when, and how without feeling overexposed or misrepresented. They described current platforms as offering rigid and limited controls that fail to account for the complexity of real social life. P16 explained that on Instagram, for example, privacy options are reduced to features like “close friends,” adding, “it’s kind of a pain to organize.” P11 similarly noted that as people share “more like their genuine real life and their experiences,” privacy needs to become “a little bit more complicated,” requiring finer-grained control over visibility.
@@ 4.5. Intentional Engagement
Participants described growing frustration with current social media platforms prioritizing passive consumption over meaningful connection. In response, they envisioned SMH as a space promoting intentional, socially rewarding interactions. This section explores themes reflecting their desire to refocus social media on authentic connection, purposeful engagement, and meaningful time use.
@@ 4.5.1. Encouraging Deliberate Navigation and Content Discovery
Participants imagined SMH encouraging deliberate, effortful discovery, replacing MSM’s passive, algorithm-driven feeds with experiences requiring users to actively choose what and who they engage with. P22 described frustration with Instagram, explaining “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage intentional, purposeful engagement when browsing content and navigating the virtual world.
@@
Participants imagined features prompting users to actively choose what they wanted to engage with before entering a space. P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces that would surface content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
Physical movement through SMH’s world was key to this intentionality. P03 explained that this kind of embodied movement makes SMH feel “more connected than regular social media,” (P03) because it requires active participation rather than passive scrolling. P02 preferred slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.” While convenient features existed, participants valued walking and exploring, turning simple actions into shared social moments.
@@
P05 believed this effortful navigation would naturally discourage superficial engagement: “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”